# A Generalized Approach to Online Convex Optimization

### v2

@@ Abstract
In this paper, we analyze the problem of online convex optimization in different settings. We show that any algorithm for online linear optimization with fully adaptive adversaries is an algorithm for online convex optimization. We also show that any such algorithm that requires full-information feedback may be transformed to an algorithm with semi-bandit feedback with comparable regret bound. We further show that algorithms that are designed for fully adaptive adversaries using deterministic semi-bandit feedback can obtain similar bounds using only stochastic semi-bandit feedback when facing oblivious adversaries. We use this to describe general meta-algorithms to convert first order algorithms to zeroth order algorithms with comparable regret bounds. Our framework allows us to analyze online optimization in various settings, such full-information feedback, bandit feedback, stochastic regret, adversarial regret and various forms of non-stationary regret, and recovers several results in the literature with a simplified proof technique.
@@ 1.1 Overview
Online optimization problems represent a class of problems where the decision-maker, referred to as an agent, aims to make sequential decisions in the face of an adversary with incomplete information Shalev-Shwartz, (2012); Hazan et al., (2016). This setting mirrors a repeated game, where the agent and the adversary engage in a strategic interplay over a finite time horizon, commonly denoted as $T$. The dynamics of this game involve the agent’s decision-making process, the adversary’s strategy, the information exchange through a query oracle.
In this paper, we present a comprehensive approach to solving online convex optimization problems, addressing scenarios where the adversary can choose a function from the class of $\mu$–strongly convex functions (or convex functions for $\mu=0$) at each time step. Our approach encompasses different feedback types, including bandit, semi-bandit and full-information feedback in each iteration. It also encompasses different types of regret, such as adversarial regret, stochastic regret, and various forms of non-stationary regret. While the problem has been studied in many special cases, this comprehensive approach sheds light on relations between different cases and the powers and limitations of some of the approaches used in different settings.
One of our key contribution lies in establishing that any semi-bandit feedback online linear (or quadratic) optimization algorithm for fully adaptive adversary is also an online convex (or strongly convex) optimization algorithm. While the above result is for semi-bandit feedback, we then show that in online convex optimization for fully adaptive adversary, semi-bandit feedback is generally enough and more information is not needed. Further, we show that algorithms that are designed for fully adaptive adversaries using deterministic semi-bandit feedback can obtain similar bounds using only stochastic semi-bandit feedback when facing oblivious adversaries. Finally, we introduce a meta algorithm, based on the variant of classical result of Flaxman et al., (2005), that converts algorithms for semi-bandit feedback, such as the ones mentioned above, to algorithms for bandit feedback.
@@
Each arrow describes a procedure or meta-algorithm to convert an algorithm from one setting into another. The first line describes the function class - linear/convex. The second lines specifies the type of the adversary - fully adaptive or oblivious. The last line starts with D or S, to denote deterministic or stochastic oracles. After that, SB is used to denotes semi-bandit, B to denote bandit and FI to denote full information feedback. Finally, the last D or S is used to denotes deterministic or stochastic algorithm.
@@ 1.2 Related works
The term “online convex optimization” was first defined in Zinkevich, (2003), where Online Gradient Descent (OGD) was proposed, while this setting was first introduced in Gordon, (1999). The Follow-The-Perturbed-Leader (FTPL) for online linear optimization was introduced by Kalai and Vempala, (2005), inspired by Follow-The-Leader algorithm originally suggested by Hannan, (1957). Follow-The-Regularized-Leader (FTRL) in the context of online convex optimization was coined in Shalev-Shwartz and Singer, (2007); Shalev-Shwartz, (2007). The equivalence of FTRL and Online Mirror Descent (OMD) was observed by Hazan and Kale, (2010). Non-stationary measures of regret have been studied in  Zinkevich, (2003); Hazan and Seshadhri, (2009); Besbes et al., (2015); Daniely et al., (2015); Zhang et al., 2018b ; Zhang et al., 2018a ; Zhao et al., (2020, 2021); Lu et al., (2023); Wang et al., (2024); Garber and Kretzu, (2022) and many other papers. We refer the reader to surveys Shalev-Shwartz, (2012) and Hazan et al., (2016) for more details.
1.3 Contributions
The key results of this paper are encapsulated in Figure 1, where each arrow represents a procedure or meta-algorithm that transforms an algorithm from one setting to another. The primary contribution is two-fold:
  • we provide a framework that simplifies analysis of meta-algorithms in online learning; and
This framework not only simplifies proof analysis for numerous existing problems in the field (see Remark 2 and Corollaries 34 and 5) but also gives a tool for easier extension of multiple results in the literature. In the following, we go over some of the main meta-algorithms that we discuss in this paper.
The first key result in this paper is Theorem 2. This result shows that any algorithm for online linear optimization with stochastic subgradient feedback results in a comparable regret with if the function class is replaced by a class of convex functions. This is the key result that allows us to go from online linear optimization to convex optimization in Figure 1 with semi-bandit feedback. We note that this result recovers the result for online convex game in Abernethy et al., (2008), where the authors analyzed online convex optimization with deterministic gradient feedback and fully adaptive adversary and showed that the min-max regret bounds of online convex optimization is equal to that of online linear optimization (or online quadratic optimization is all convex functions are $\mu$-strongly convex).
The second key result, given in Theorem 3, is that for any full-information algorithm for online convex optimization with fully adaptive adversary, there is a corresponding algorithm with regret bound of the same order which only requires semi-bandit feedback. This allows us to get results for semi-bandit feedback from that of full information. This combined with Theorem 2 allows us to go from any full-information algorithm for online linear optimization with fully adaptive adversary to an algorithm for online convex optimization with semi-bandit feedback and fully adaptive adversary. This combination has been marked as Theorem 3 in Figure 1.
We note that the above results explained in Theorems 14 assume that the feedback oracle is deterministic. In the next key result, given in Theorem 5, we relax this assumption. We show that any online optimization algorithm for fully adaptive adversaries using deterministic semi-bandit feedback obtains similar regret bounds when facing an oblivious adversary using only stochastic semi-bandit feedback. Thus, this result allows us to get the results for online convex optimization with stochastic semi-bandit feedback from the results with deterministic semi-bandit feedback, as can be seen in Fig. 1. This result gives Theorem 6.5 in Hazan et al., (2016) as a special case. Further, this allows us to have a result that generalizes both adversarial regret and stochastic regret, as they are commonly defined in the literature.
Finally, we note that the above discussion was mostly focused on full information and semi-bandit feedback settings. However, in many practical setups, zeroth order or bandit feedback is important. In the final key result, given in Theorems 6-7, we use the smoothing trick Flaxman et al., (2005); Pedramfar et al., (2023), in order to estimate the gradient from the value oracle. Using this estimated gradient, the results for zeroth order and bandit feedback are derived from that for first order and semi-bandit feedback, respectively. These results apply to any optimization problem, not necessarily convex. This approach recovers the results in Flaxman et al., (2005) with the base algorithm of OGD, while extends to say that the approach can be applied directly on any base algorithm. In particular, this approach also matches the SOTA dynamic and adaptive regret bounds for online convex optimization with bandit feedback Zhao et al., (2021); Garber and Kretzu, (2022) with a simplified proof.
Our work sheds light on the relation between different problems, algorithms and techniques used in online optimization and provides general meta-algorithms that allow conversion between different problem settings. For instance, we can use the results for a deterministic algorithm for online linear optimization with oblivious adversaries and deterministic semi-bandit feedback to that for a stochastic algorithm for online convex optimization with oblivious adversaries and stochastic bandit feedback using the different arrows as in Fig. 1. The key technical novelty is the recognition that many results, previously understood only in very limited contexts, actually apply more broadly. This broader applicability not only simplifies the analysis of numerous existing results within the field but also enables the derivation of multiple new results. For instance, results in Abernethy et al., (2008); Hazan et al., (2016); Flaxman et al., (2005); Zhao et al., (2021) as mentioned above can be demonstrated as corollaries of these more generalized results.
@@ 2 Background and Notations
For a set ${\mathcal{D}}\subseteq{\mathbb{R}}^{d}$, we define its affine hull $\operatorname{aff}({\mathcal{D}})$ to be the set of $\alpha{\mathbf{x}}+(1-\alpha){\mathbf{y}}$ for all ${\mathbf{x}},{\mathbf{y}}$ in ${\mathcal{K}}$ and $\alpha\in{\mathbb{R}}$. The relative interior of ${\mathcal{D}}$ is defined as
@@
A function class is simply a set of real-valued functions. Given a set ${\mathcal{D}}$, a function class over ${\mathcal{D}}$ is simply a subset of all real-valued functions on ${\mathcal{D}}$. Given a real number $\mu\geq 0$ and a set ${\mathcal{D}}$, we use ${\mathbf{Q}}_{\mu}$ to denote the class of functions of the form $q_{{\mathbf{x}},{\mathbf{o}}}({\mathbf{y}})=\langle{\mathbf{o}},{\mathbf{y}}-{\mathbf{x}}\rangle+\frac{\mu}{2}\|{\mathbf{y}}-{\mathbf{x}}\|^{2}$, where ${\mathbf{o}}\in{\mathbb{R}}^{d}$ and ${\mathbf{x}}\in{\mathcal{D}}$. A set ${\mathcal{K}}\subseteq{\mathbb{R}}^{d}$ is called a convex set if for all ${\mathbf{x}},{\mathbf{y}}\in{\mathcal{K}}$ and $\alpha\in[0,1]$, we have $\alpha{\mathbf{x}}+(1-\alpha){\mathbf{y}}\in{\mathcal{K}}$. Given a convex set ${\mathcal{K}}$, a function $f:{\mathcal{K}}\to{\mathbb{R}}$ is called convex if, for all ${\mathbf{x}},{\mathbf{y}}\in{\mathcal{K}}$ and $\alpha\in[0,1]$, we have $f(\alpha{\mathbf{x}}+(1-\alpha){\mathbf{y}})\leq\alpha f({\mathbf{x}})+(1-\alpha)f({\mathbf{y}})$. All convex functions are continuous on any point in the relative interior of their domains. In this work, we will only focus on continuous functions. If ${\mathbf{x}}\in\operatorname{relint}({\mathcal{K}})$ and $f$ is convex and is differentiable at ${\mathbf{x}}$, then we have $f({\mathbf{y}})-f({\mathbf{x}})\geq\langle\nabla f({\mathbf{x}}),{\mathbf{y}}-{\mathbf{x}}\rangle$, for all ${\mathbf{y}}\in{\mathcal{K}}$. A vector ${\mathbf{o}}\in{\mathbb{R}}^{d}$ is called a subgradient of $f$ at ${\mathbf{x}}$ if $f({\mathbf{y}})-f({\mathbf{x}})\geq\langle{\mathbf{o}},{\mathbf{y}}-{\mathbf{x}}\rangle$, for all ${\mathbf{y}}\in{\mathcal{K}}$. Given a convex set ${\mathcal{K}}$, a function $f:{\mathcal{K}}\to{\mathbb{R}}$ is convex if and only if it has a subgradient at all points ${\mathbf{x}}\in{\mathcal{K}}$. More generally, given $\mu\geq 0$, we say a vector ${\mathbf{o}}\in{\mathbb{R}}^{d}$ is a $\mu$-subgradient of $f$ at ${\mathbf{x}}$ if $f({\mathbf{y}})-f({\mathbf{x}})\geq\langle{\mathbf{o}},{\mathbf{y}}-{\mathbf{x}}\rangle+\frac{\mu}{2}\|{\mathbf{y}}-{\mathbf{x}}\|^{2}$, for all ${\mathbf{y}}\in{\mathcal{K}}$. Given a convex set ${\mathcal{K}}$, a function $f:{\mathcal{K}}\to{\mathbb{R}}$ is $\mu$-strongly convex if and only if it has a $\mu$-subgradient at all points ${\mathbf{x}}\in{\mathcal{K}}$.
@@ 3 Problem Setup
Online optimization problems can be formalized as a repeated game between an agent and an adversary. The game lasts for $T$ rounds on a convex domain ${\mathcal{K}}$ where $T$ and ${\mathcal{K}}$ are known to both players. In $t$-th round, the agent chooses an action ${\mathbf{x}}_{t}$ from an action set ${\mathcal{K}}\subseteq{\mathbb{R}}^{d}$, then the adversary chooses a loss function $f_{t}\in{\mathbf{F}}$ and a query oracle for the function $f_{t}$. Then, for $1\leq i\leq k_{t}$, the agent chooses a points ${\mathbf{y}}_{t,i}$ and receives the output of the query oracle.
@@
Static adversarial regret or simply adversarial regret corresponds to $a=1$, $b=T$ and ${\mathcal{U}}={\mathcal{K}}_{\star}^{T}:=\{({\mathbf{x}},\cdots,{\mathbf{x}})\mid{\mathbf{x}}\in{\mathcal{K}}\}$. When $a=1$, $b=T$ and ${\mathcal{U}}$ contains only a single element then it is referred to as the dynamic regret Zinkevich, (2003); Zhang et al., 2018a . If instead ${\mathcal{U}}={\mathcal{K}}^{T}$, then it is referred to as restricted dynamic regret Besbes et al., (2015). Adaptive regret, is defined as $\max_{1\leq a\leq b\leq T}{\mathcal{R}}_{\operatorname{Adv}}^{{\mathcal{A}}}({\mathcal{K}}_{\star}^{T})[a,b]$ Hazan and Seshadhri, (2009). We drop $a$, $b$ and ${\mathcal{U}}$ when the statement is independent of their value or their value is clear from the context.

### v3

# A Unified Framework for Analyzing Meta-algorithms in Online Convex Optimization

@@ Abstract
In this paper, we analyze the problem of online convex optimization in different settings, including different feedback types (full-information/semi-bandit/bandit/etc) in either stochastic or non-stochastic setting and different notions of regret (static adversarial regret/dynamic regret/adaptive regret). This is done through a framework which allows us to systematically propose and analyze meta-algorithms for the various settings described above. We show that any algorithm for online linear optimization with fully adaptive adversaries is an algorithm for online convex optimization. We also show that any such algorithm that requires full-information feedback may be transformed to an algorithm with semi-bandit feedback with comparable regret bound. We further show that algorithms that are designed for fully adaptive adversaries using deterministic semi-bandit feedback can obtain similar bounds using only stochastic semi-bandit feedback when facing oblivious adversaries. We use this to describe general meta-algorithms to convert first order algorithms to zeroth order algorithms with comparable regret bounds. Our framework allows us to analyze online optimization in various settings, recovers several results in the literature with a simplified proof technique, and provides new results.
@@ 1.1 Overview
Online optimization problems represent a class of problems where the decision-maker, referred to as an agent, aims to make sequential decisions in the face of an adversary with incomplete information (?, ?). This setting mirrors a repeated game, where the agent and the adversary engage in a strategic interplay over a finite time horizon, commonly denoted as $T$. The dynamics of this game involve the agent’s decision-making process, the adversary’s strategy, and the information exchange through a query oracle.
In this paper, we present a comprehensive approach to solve online convex optimization problems, addressing scenarios where the adversary can choose a function from the class of $\mu$–strongly convex functions (or convex functions for $\mu=0$) at each time step. Our approach encompasses different feedback types, including bandit, semi-bandit and full-information feedback in each iteration. It also encompasses different types of regret, such as adversarial regret, stochastic regret, and various forms of non-stationary regret. While the problem has been studied in many special cases, this comprehensive approach sheds light on relations between different cases and the powers and limitations of some of the approaches used in different settings.
One of our key contribution lies in establishing that any semi-bandit feedback online linear (or quadratic) optimization algorithm for fully adaptive adversary is also an online convex (or strongly convex) optimization algorithm. While the above result is for semi-bandit feedback, we then show that in online convex optimization for fully adaptive adversary, semi-bandit feedback is generally enough and more information is not needed. Further, we show that algorithms that are designed for fully adaptive adversaries using deterministic semi-bandit feedback can obtain similar bounds using only stochastic semi-bandit feedback when facing oblivious adversaries. Finally, we introduce meta algorithms, based on the variant of classical result of (?, ?), that convert algorithms for semi-bandit feedback, such as the ones mentioned above, to algorithms for bandit or zeroth-order feedback. In addition to recovering many results in the literature with simplified proofs, we give new static regret guarantees for online strongly convex optimization and adaptive regret guarantees for convex optimization with deterministic zeroth order feedback.
@@
Each arrow describes a procedure or meta-algorithm to convert an algorithm from one setting into another. The first line describes the function class - linear/convex. The second lines specifies the type of the adversary - fully adaptive or oblivious. The last line starts with D or S, to denote deterministic or stochastic oracles. After that, SB is used to denotes semi-bandit, B to denote bandit and FI to denote full information feedback. Finally, the last D or S, if exists, is used to denotes deterministic or stochastic algorithm.
1.2 Contributions
The key results of this paper are encapsulated in Figure 1, where each arrow represents a procedure or meta-algorithm that transforms an algorithm from one setting to another. The primary contribution is three-fold:
  • (I)
    We provide a framework that simplifies analysis of meta-algorithms in online learning.
  • (II)
  • (III)
    In particular, we convert several insights in the literature into precise theorems, for example the idea of using 1-point gradient estimator in online optimization is formalized in Meta-algorithms 2 and 3 and Theorems 6 and 7. Similarly, the insight that “knowing the value of each loss function at two points is almost as useful as knowing the value of each function everywhere” (See (?)) is formalized in Meta-algorithm 4 and Theorem 8.
This framework not only simplifies proof analysis for numerous existing problems in the field (see Section 11) but also provides a tool for easier extension of multiple results in the literature. In the following, we go over some of the main meta-algorithms that we discuss in this paper.
(i) The first key result in this paper is Theorem 2. This result shows that any algorithm for online linear optimization with stochastic subgradient feedback results in a comparable regret with if the function class is replaced by a class of convex functions. This is the key result that allows us to go from online linear optimization to convex optimization in Figure 1 with semi-bandit feedback. We note that this result recovers the result for online convex game in (?), where the authors analyzed online convex optimization with deterministic gradient feedback and fully adaptive adversary and showed that the min-max regret bounds of online convex optimization is equal to that of online linear optimization (or online quadratic optimization is all convex functions are $\mu$-strongly convex).
(ii) The second key result, given in Theorem 3, is that for any full-information algorithm for online convex optimization with fully adaptive adversary, there is a corresponding algorithm with regret bound of the same order which only requires semi-bandit feedback. This allows us to get results for semi-bandit feedback from that of full information. This combined with Theorem 2 allows us to go from any full-information algorithm for online linear optimization with fully adaptive adversary to an algorithm for online convex optimization with semi-bandit feedback and fully adaptive adversary. This combination has been marked as Theorem 3 in Figure 1.
(iii) We note that the above results assume that the feedback oracle is deterministic. In the next key result, given in Theorem 5, we relax this assumption. We show that any online optimization algorithm for fully adaptive adversaries using deterministic semi-bandit feedback obtains similar regret bounds when facing an oblivious adversary using only stochastic semi-bandit feedback. Thus, this result allows us to get the results for online convex optimization with stochastic semi-bandit feedback from the results with deterministic semi-bandit feedback, as can be seen in Fig. 1. This result gives Theorem 6.5 in (?) as a special case. Further, this allows us to have a result that generalizes both adversarial regret and stochastic regret, as they are commonly defined in the literature.
(iv) The above discussion was mostly focused on full information and semi-bandit feedback settings. However, in many practical setups, we only have access to zeroth order or bandit feedback. In the next key result, given in Theorems 6-7, we use the smoothing trick (?, ?, ?), in order to estimate the gradient from the value oracle. Using this estimated gradient, the results for zeroth order and bandit feedback are derived from that for first order and semi-bandit feedback, respectively. These results apply to any optimization problem, not necessarily convex. This approach recovers the results in (?) with the base algorithm of Online Gradient Descent ($\mathtt{OGD}$). In particular, this approach also matches the SOTA dynamic and adaptive regret bounds for online convex optimization with bandit feedback (?, ?, ?) with a simplified proof.
(v) In the above results, we assumed that we only have access to an unbiased estimate of the zeroth order (or bandit) feedback. In (?), the setting of multi-point bandit feedback is introduced. Even though their setting and notion of regret is different from our work, one insight is relevant. Their analysis suggests that, in some cases, if we have access to exact values of the functions, knowing the value of each loss function at two points is almost as useful as knowing the value of each function everywhere. We formalize this insight by designing a meta-algorithm based on the two-point gradient estimator (?, ?) that allows us to convert algorithms for stochastic first order feedback into algorithms for deterministic zeroth order feedback with the same order of regret bound (Theorem 8). Note that this result holds for convex and non-convex optimization as it only relies on the function class being Lipschitz. This result is shown to recover the result for dynamic regret for Improved Ader (?), while achieving two new results for adaptive regret for convex function of $O(\sqrt{T})$ and static regret for strongly convex function of $O(\log T)$.
Our work sheds light on the relation between different problems, algorithms and techniques used in online optimization and provides general meta-algorithms that allow conversion between different problem settings. For instance, we can use the results for a deterministic algorithm for online linear optimization with oblivious adversaries and deterministic semi-bandit feedback to that for a stochastic algorithm for online convex optimization with oblivious adversaries and stochastic bandit feedback using the different arrows as in Fig. 1. The key technical novelty is the recognition that many results, previously understood only in very limited contexts, actually apply more broadly. This broader applicability not only simplifies the analysis of numerous existing results within the field but also enables the derivation of multiple new results. For instance, results in (?, ?, ?, ?, ?, ?) as mentioned above can be demonstrated as corollaries of these more generalized results. Further, we provide new results, which shows that adaptive regret for convex function is $O(\sqrt{T})$ and static regret for strongly convex function is $O(\log T)$.
@@ 1.3 Related works
The term “online convex optimization” was first defined in (?), where Online Gradient Descent ($\mathtt{OGD}$) was proposed, while this setting was first introduced in (?). The Follow-The-Perturbed-Leader ($\mathtt{FTPL}$) for online linear optimization was introduced by (?), inspired by Follow-The-Leader algorithm originally suggested by (?). Follow-The-Regularized-Leader ($\mathtt{FTRL}$) in the context of online convex optimization was coined in (?, ?). The equivalence of $\mathtt{FTRL}$ and Online Mirror Descent ($\mathtt{OMD}$) was observed by (?). Non-stationary measures of regret have been studied in  (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?) and many other papers. We refer the reader to surveys (?) and (?) for more details.
@@ 2 Background and Notations
For a set ${\mathcal{D}}\subseteq{\mathbb{R}}^{d}$, we define its affine hull $\operatorname{aff}({\mathcal{D}})$ to be the set of $\alpha{\mathbf{x}}+(1-\alpha){\mathbf{y}}$ for all ${\mathbf{x}},{\mathbf{y}}$ in ${\mathcal{D}}$ and $\alpha\in{\mathbb{R}}$. The relative interior of ${\mathcal{D}}$ is defined as
@@
We use ${\mathbb{B}}_{r}({\mathcal{D}})$ to denote the union of balls of radius $r$ centered at ${\mathcal{D}}$. A function class is simply a set of real-valued functions. Given a set ${\mathcal{D}}$, a function class over ${\mathcal{D}}$ is simply a subset of all real-valued functions on ${\mathcal{D}}$. Given a real number $\mu\geq 0$ and a set ${\mathcal{D}}$, we use ${\mathbf{Q}}_{\mu}$ to denote the class of functions of the form $q_{{\mathbf{x}},{\mathbf{o}}}({\mathbf{y}})=\langle{\mathbf{o}},{\mathbf{y}}-{\mathbf{x}}\rangle+\frac{\mu}{2}\|{\mathbf{y}}-{\mathbf{x}}\|^{2}$, where ${\mathbf{o}}\in{\mathbb{R}}^{d}$ and ${\mathbf{x}}\in{\mathcal{D}}$. A set ${\mathcal{K}}\subseteq{\mathbb{R}}^{d}$ is called a convex set if for all ${\mathbf{x}},{\mathbf{y}}\in{\mathcal{K}}$ and $\alpha\in[0,1]$, we have $\alpha{\mathbf{x}}+(1-\alpha){\mathbf{y}}\in{\mathcal{K}}$. Given a convex set ${\mathcal{K}}$, a function $f:{\mathcal{K}}\to{\mathbb{R}}$ is called convex if, for all ${\mathbf{x}},{\mathbf{y}}\in{\mathcal{K}}$ and $\alpha\in[0,1]$, we have $f(\alpha{\mathbf{x}}+(1-\alpha){\mathbf{y}})\leq\alpha f({\mathbf{x}})+(1-\alpha)f({\mathbf{y}})$. All convex functions are continuous on any point in the relative interior of their domains, but not necessarily at their relative boundary. In this work, we will only focus on continuous functions. If ${\mathbf{x}}\in\operatorname{relint}({\mathcal{K}})$ and $f$ is convex and is differentiable at ${\mathbf{x}}$, then we have $f({\mathbf{y}})-f({\mathbf{x}})\geq\langle\nabla f({\mathbf{x}}),{\mathbf{y}}-{\mathbf{x}}\rangle$, for all ${\mathbf{y}}\in{\mathcal{K}}$. A vector ${\mathbf{o}}\in{\mathbb{R}}^{d}$ is called a subgradient of $f$ at ${\mathbf{x}}$ if $f({\mathbf{y}})-f({\mathbf{x}})\geq\langle{\mathbf{o}},{\mathbf{y}}-{\mathbf{x}}\rangle$, for all ${\mathbf{y}}\in{\mathcal{K}}$. Given a convex set ${\mathcal{K}}$, a function $f:{\mathcal{K}}\to{\mathbb{R}}$ is convex if and only if it has a subgradient at all points ${\mathbf{x}}\in{\mathcal{K}}$. More generally, given $\mu\geq 0$, we say a vector ${\mathbf{o}}\in{\mathbb{R}}^{d}$ is a $\mu$-subgradient of $f$ at ${\mathbf{x}}$ if $f({\mathbf{y}})-f({\mathbf{x}})\geq\langle{\mathbf{o}},{\mathbf{y}}-{\mathbf{x}}\rangle+\frac{\mu}{2}\|{\mathbf{y}}-{\mathbf{x}}\|^{2}$, for all ${\mathbf{y}}\in{\mathcal{K}}$. Given a convex set ${\mathcal{K}}$, a function $f:{\mathcal{K}}\to{\mathbb{R}}$ is $\mu$-strongly convex if and only if it has a $\mu$-subgradient at all points ${\mathbf{x}}\in{\mathcal{K}}$. We use $\nabla^{*}f$ to denote the set of $\mu$-subgradients of $f$.
@@ 3 Problem Setup
Online optimization problems can be formalized as a repeated game between an agent and an adversary. The game lasts for $T$ rounds on a convex domain ${\mathcal{K}}$ where $T$ and ${\mathcal{K}}$ are known to both players. In $t$-th round, the agent chooses an action ${\mathbf{x}}_{t}$ from an action set ${\mathcal{K}}\subseteq{\mathbb{R}}^{d}$, then the adversary chooses a loss function $f_{t}\in{\mathbf{F}}$ and a query oracle for the function $f_{t}$. Then, for $1\leq i\leq k_{t}$, the agent chooses a points ${\mathbf{y}}_{t,i}$ and receives the output of the query oracle. Here $k_{t}$ denotes the total number of queries made by the agent at time-step $t$, which may or may not be known in advance.
@@
Static adversarial regret or simply adversarial regret corresponds to $a=1$, $b=T$ and ${\mathcal{U}}={\mathcal{K}}_{\star}^{T}:=\{({\mathbf{x}},\cdots,{\mathbf{x}})\mid{\mathbf{x}}\in{\mathcal{K}}\}$. When $a=1$, $b=T$ and ${\mathcal{U}}$ contains only a single element then it is referred to as the dynamic regret (?, ?). If instead ${\mathcal{U}}={\mathcal{K}}^{T}$, then it is referred to as restricted dynamic regret (?). Adaptive regret, is defined as $\max_{1\leq a\leq b\leq T}{\mathcal{R}}_{\operatorname{Adv}}^{{\mathcal{A}}}({\mathcal{K}}_{\star}^{T})[a,b]$ (?). We drop $a$, $b$ and ${\mathcal{U}}$ when the statement is independent of their value or their value is clear from the context.